a magnetic field is given by the expression b=axzi + byzj + ck. what is the relationship between a and b


1
Expert's answer
2021-05-17T18:55:12-0400

Given magnetic field

B=axi+byj+ck

We know that

Maxwell second equation

"\nabla.B=0" "\rightarrow(1)"

equation (1) put value

"(i\frac{d}{dx}+j\frac{d}{dy}+k\frac{d}{dz}).(axzi+byzj+ck)=0"

=az+bz=0

z(a+b)=0

a+b=0